1) Refer to the output table below. Assume 2014 is the base year.

Year Good 1 Price Good 1 Output Good 2 Price Good 2 Output
2014 $0.50 100 $12.00 50
2015 0.70 102 13.00 54
2016 0.75 105 14.00 58
2017 1.00 120 15.00 60

Instructions: Round your answer 2 decimal places.

Real GDP in 2014 is $   .

Nominal GDP in 2017 is $ ______
Answer #1

Ans) real GDP in 2014 = 0.50* 100 + 12*50

= 50 + 600 = 650

Nominal GDP in 2017 = 1*120 + 15*60 = 1020